# SAQR-QC: A Logic for Scalable but Approximate Quantitative Reasoning about Quantum Circuits

### v1

@@ 2.2. Quantum States
A quantum state describes the state of a quantum system. For a single qubit, the state |$\ket{\psi}$ belongs to a two-dimensional Hilbert space and can exist as a superposition of the basis states |$\ket{0}$ and |$\ket{1}$. For an $n$-qubit system, the state resides in a $2^{n}$-dimensional Hilbert space and can exhibit both complex superpositions and entanglement among qubits.
Quantum systems may also be in mixed states, represented by a density matrix $\rho$, which generalizes pure states to account for probabilistic mixtures—capturing both classical and quantum uncertainty. Quantum operations, whether on pure or mixed states, include unitary transformations (which preserve total probability) and measurements, which probabilistically collapse the system to a classical outcome based on the state’s amplitudes.
@@ 2.3. Reduced Density Matrices
Let $\mathbb{C}^{d_{1}}$ and $\mathbb{C}^{d_{2}}$ be the Hilbert spaces of two quantum systems. The state space of the composite system is modeled by the tensor product $\mathbb{C}^{d_{1}}\otimes\mathbb{C}^{d_{2}}$, and analyzing subsystems requires the notion of the partial trace. Formally, the partial trace over $\mathbb{C}^{d_{1}}$, denoted ${\mathrm{Tr}}_{1}(\cdot)$, maps operators on $\mathbb{C}^{d_{1}}\otimes\mathbb{C}^{d_{2}}$ to operators on $\mathbb{C}^{d_{2}}$, and is defined by
@@
for all |$\ket{\varphi_{1}},\ket{\psi_{1}}\in\mathbb{C}^{d_{1}}$ and |$\ket{\varphi_{2}},\ket{\psi_{2}}\in\mathbb{C}^{d_{2}}$, and extended linearly. Similarly, ${\mathrm{Tr}}_{2}(\cdot)$ denotes the partial trace over $\mathbb{C}^{d_{2}}$. For a composite system with density matrix $\rho\in\mathbb{C}^{d_{1}}\otimes\mathbb{C}^{d_{2}}$, the reduced states of the first and second subsystems are given by ${\mathrm{Tr}}_{2}(\rho)$ and ${\mathrm{Tr}}_{1}(\rho)$, respectively.
The notion of partial trace extends naturally to $n$-partite systems: for a subset $s\subseteq[n]$, the reduced density matrix of the subsystem indexed by $s$ is given by
@@
where ${\mathrm{Tr}}_{[n]\setminus s}$ denotes tracing out all qubits not in $s$. The partial-trace operation preserves positive semi-definiteness (NI11, ).
Reduced density matrices play a central role in the analysis of multipartite quantum systems. Many important properties of a quantum system depend solely on the reduced density matrices of its subsystems. This principle holds in quantum computation as well. For instance, the success probability of many quantum algorithms—such as the well-known HHL algorithm (HHL08, )—depends solely on the reduced density matrix of a specific qubit (the “signal” qubit). In such algorithms, the signal qubit is measured, and the computation is considered successful if the outcome is |$\ket{1}$. Importantly, the probability of this outcome is fully determined by the reduced density matrix of the signal qubit, regardless of the entanglement or structure of the global quantum state.
@@ 2.4. Unitary Operations
Unitary operations are fundamental transformations in quantum mechanics; they preserve the norm of a quantum state and are represented by a unitary matrix. A unitary matrix $U$ satisfies $U^{\dagger}U=I$, where $U^{\dagger}$ denotes the conjugate transpose of $U$ and $I$ is the identity matrix. These operations are crucial for manipulating quantum states and implementing quantum algorithms. For a pure state |$\ket{\psi}$, applying a unitary operator $U$ transforms |$\ket{\psi}$ to $U\ket{\psi}$. For a density operator $\rho$, the transformation is $\rho\mapsto U\rho U^{\dagger}$.
Commonly used single-qubit operators include the Hadamard gate $H$, the $T$ gate, the family of gates $\set{R_{m}\mid m\in\mathbb{N}}$ and the Pauli gates $I$, $X$, $Y$, and $Z$. Commonly used two-qubit gates include the SWAP operation SWAP and the controlled-NOT operation CNOT.
@@ 2.5. Observables
Quantum observables are physical quantities in a quantum system that can be measured, such as position, momentum, energy, and spin. Mathematically, observables are represented by Hermitian (self-adjoint) operators on a Hilbert space, satisfying $O^{\dagger}=O$, which ensures that measurement outcomes are real. In certain contexts—such as Quantum Hoare Logic (QHL)—it is common to further restrict observables to the range $0\leq O\leq I$, where the inequality is understood in the Löwner partial order (§ 2.1), meaning that both $O$ and $I-O$ are positive semidefinite. These bounded observables often serve as predicates or quantum effects, capturing partial truth values within the quantum program-verification framework.
@@ 2.6. Quantum Circuits and Semantics
In this work, we focus on quantum programs—represented as quantum circuits—operating on a fixed number of qubits, says $n$. A quantum program is composed of a sequence of unitary instructions $U_{F_{1}},\dots,U_{F_{|p|}}$. Each gate $U_{F_{\ell}}$ operates on a subset of qubits $F_{\ell}\subseteq[n]$. The initial state is taken to be |$\ket{0}^{\otimes n}=\ket{0^{n}}$, and the meaning (or semantics) of the program is given by the matrix product
$U_{F_{|p|}}\cdots U_{F_{1}}\ket{0^{n}}.$
To interpret $U_{F_{\ell}}$ as an $n$-qubit unitary, we embed it into the full register by tensoring with identity operators. If $F_{\ell}=\{i\}$, then the corresponding lifted unitary is
$U\otimes I_{[n]\setminus\{i\}}:=(\otimes_{k>i}I)\otimes U\otimes(\otimes_{0\leq j<i}I),$
where $U$ acts nontrivially on qubit $i$, and $I$ is the identity matrix on one qubit. We will explicitly subscript identity matrices to indicate which qubits they act upon.
For a two-qubit unitary $U_{F_{\ell}}$ acting on qubits $F_{\ell}=\{i,j\}$, we similarly interpret it as
$U\otimes I_{[n]\setminus\{i,j\}},$
with the appropriate placement determined by the positions of $i$ and $j$ in the register.
To simplify the presentation and reduce notational overhead, we formulate the semantics of quantum circuits assuming 2-qubit gates. Nevertheless, all of our methods naturally extend to circuits composed of gates acting on up to $m$ qubits, for any constant $m$.
@@
We also define the dual action on observables as:
$\llbracket C\rrbracket^{*}(A):=U_{C}^{\dagger}AU_{C},$
for any observable $A$.
@@ 2.7. Projections
An orthogonal projection matrix $P$ satisfies $P=P^{\dagger}=P^{2}$, which a stricter condition than the classical $P=P^{2}$. For short, we refer to such matrices as “projections.” For example, $|00\rangle\langle 00|+|11\rangle\langle 11|$ is a rank-2 projection that projects any 4-dimensional vector onto a 2-dimensional subspace.
Each projection $P$ corresponds to a unique subspace $S_{P}=\{v\mid Pv=v\}$, and we use the terms “projections” and “subspaces” interchangeably. This correspondence establishes a partial order: for projections $P$ and $Q$, we have $P\subseteq Q$ iff $S_{P}\subseteq S_{Q}$.
Projections are positive semi-definite. The support of a positive semi-definite matrix $A$, $\operatorname{supp}(A)$, is the subspace spanned by eigenvectors with nonzero eigenvalues. According to Birkhoff and von Neumann (BvN36, ), a density matrix $\rho$ satisfies a projection $P$, denoted by $\rho\vDash P$, if $\operatorname{supp}(\rho)\subseteq P$. This property is equivalent to $P\rho=\rho$.
@@ 2.8. Lemmas
The following lemmas capture algebraic relationships among these operations. These lemmas are instrumental in establishing the correctness of our SAQR-QC framework in a purely algebraic style. In what follows, we assume $s\subseteq[n]$; that $P$ is a projection operator on $n$-qubit systems; that $A,B$ is a positive semidefinite matrix on an $n$-qubit space, $E$ be a matrix; and that $\rho$ is a quantum state.
@@
Let $\rho$ be the density matrix of an $n$-qubit system, and let $s\subseteq[n]$. Then for any observable $A_{s}$ acting on subsystem $s$,
@@
If $A\leq B$, then ${\mathrm{Tr}}(A\rho)\leq{\mathrm{Tr}}(B\rho)$ for any density operator $\rho$. In particular, ${\mathrm{Tr}}(A)\leq{\mathrm{Tr}}(B)$.
@@
The proof of 2.1 is given in Appendix A. The other lemmas can be proven using the definition of trace, partial trace, and support.
3. Qualitative Predicates for Local Reasoning: Quantum Abstract Interpretation (YP21, )
Qualitative predicates serve as logical assertions about quantum states, such as whether a subsystem lies in a given subspace. When expressed as tuples of local projectors, they enable scalable local reasoning, inferring global behavior from partial system views. Quantum Abstract Interpretation (QAI) (YP21, ) systematically propagates such predicates through quantum circuits using projectors on small subsystems, avoiding the exponential cost of full-state analysis.
This section reviews local projective predicates, formalizes their semantics, and describes their transformation under unitaries. Soundness is established via support-based semantics and partial trace, forming a foundation for abstract reasoning in quantum programs.
@@ 3. Qualitative Predicates for Local Reasoning: Quantum Abstract Interpretation (YP21, )
An $n$-qubit quantum state $\rho$ satisfies a projective predicate $\mathcal{P}=(P_{s_{1}},\cdots,P_{s_{m}})$, denoted by
@@
if for all $1\leq i\leq m$, we have $P_{s_{i}}\rho_{s_{i}}=\rho_{s_{i}}$, i.e., $\rho_{s_{i}}\vDash P_{s_{i}}$. Equivalently, this property holds if $\rho\vDash\gamma(\mathcal{P})$, where
@@
Let $P=\mathrm{span}\{\ket{a_{1}a_{2}\cdots a_{n}},\ket{b_{1}b_{2}\cdots b_{n}}\}$, where the product states |$\ket{a_{i}}$ and |$\ket{b_{i}}$ are not parallel for every $i\in[n]$. Then
$P=\gamma(\mathcal{P}),$
where $\mathcal{P}=(P_{1,2},\ldots,P_{n-1,n})$ and each $P_{i,i+1}=\mathrm{span}\{\ket{a_{i}a_{i+1}},\ket{b_{i}b_{i+1}}\}.$
@@ 4. Correctness Formulas and the Logical System SAQR-QC
In this section, we first motivate the use of local observables in the definition of quantum predicates. We then apply this idea to Quantum Hoare Logic (QHL) as an initial attempt to define a corresponding judgment. Through a simple example, we demonstrate that this approach can lead to a significant loss of precision in reasoning. To address this limitation, we subsequently integrate the QAI technique into the framework, leading to the formal definition of the judgments used in SAQR-QC and a systematic presentation of the inference rules of SAQR-QC.
4.1. Motivation.
In the analysis and verification of quantum programs, reasoning about the full quantum state quickly becomes infeasible due to the exponential growth of the state space with the number of qubits. To address this challenge, we shift our focus to reduced density matrices, which capture the behavior of local subsystems. Notably, a tuple of reduced density matrices—each defined on a constant number of qubits—requires only a linear number of parameters to describe. This property stands in stark contrast to the exponential size of the global state and provides a practical pathway to tractable reasoning.
Let $(\rho_{s_{1}},\dots,\rho_{s_{m}})$ denote a tuple of reduced density matrices, each corresponding to a small subset $s_{i}\subseteq[n]$ of the full system. Inspired by quantum Hoare logic (QHL), where a global positive semi-definite matrix $A$ is used as a predicate to track changes in the global state $\rho$, we generalize this approach by employing linear functionals—namely, local observables—to track the evolution of each $\rho_{s_{i}}$. This approach allows us to perform quantitative reasoning over the program’s behavior through its low-dimensional marginals.
Local observables, which act nontrivially on only a few qubits, serve as efficient probes for monitoring how the reduced density matrices evolve during program execution. Because observable expectation values determine measurable quantities like success probabilities, fidelities, and entanglement, they form a natural bridge between the semantics of quantum programs and their operational outcomes.
By tracing how these observables evolve under unitary transformations and partial traces, we establish a scalable and compositional framework for reasoning about quantum programs. This approach enables precise quantitative analysis without requiring reconstruction of the global state, laying the groundwork for scalable verification of large-scale quantum computations through their locally observable structure.
@@ 4.2. Quantitative Local Reasoning via Generalized Predicates
This subsection defines predicates for quantitative local reasoning by generalizing from local projections to local observables. While qualitative reasoning (as in (YP21, )) relies on tuples of projections to characterize logical properties of subsystems, our approach extends this framework to encompass a broader class of observables—specifically, positive semidefinite operators bounded by the identity.
These generalized predicates enable us to track quantitative aspects of quantum programs, such as probabilities and expectation values, by monitoring the evolution of reduced density matrices over small, constant-size subsystems. This generalization forms the foundation for scalable, compositional reasoning about quantum computations through low-dimensional, information-preserving abstractions.
@@
Given an $n$-qubit Hilbert space, an integer $m$, and an $m$-tuple $S=(s_{1},\cdots,s_{m})$ with each $s_{i}\subseteq[n]$ indicating a subset of qubits, the set of predicates is defined as:
[…]
where each $A_{s_{i}}$ is a local observable acting nontrivially only on subsystem $s_{i}$. We usually use $\mathscr{A}$ (or $\mathscr{B}$, $\mathscr{D}$) to denote predicates. The integer $m$ is called the size of the predicate.
@@
$M_{\mathscr{A}}=\sum_{i=1}^{m}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}.$
@@
The domain of a predicate $(A_{s_{1}},\cdots,A_{s_{m}})$, denoted by $\mathrm{dom}(A_{s_{1}},\cdots,A_{s_{m}})$, is $(s_{1},\cdots,s_{m})$.
We will see that the above definition of predicates serves to track changes in a specified tuple of reduced density matrices by means of linear functionals, facilitating quantitative reasoning in a scalable and compositional manner. By Lemma 2.1, the trace ${\mathrm{Tr}}(M_{\mathscr{A}}\rho)$ depends only on the reduced density matrices $\rho_{s_{i}}$ of $\rho$:
(6) ${{\mathrm{Tr}}}(M_{\mathscr{A}}\rho)={\mathrm{Tr}}\left(\sum_{i=1}^{m}(A_{s_{i}}\otimes I_{[n]\setminus s_{i}})\rho\right)=\sum_{i=1}^{m}{\mathrm{Tr}}[(A_{s_{i}}\otimes I_{[n]\setminus s_{i}})\rho]=\sum_{i=1}^{m}{\mathrm{Tr}}(A_{s_{i}}\rho_{s_{i}}).$
@@
Given two predicates $(A_{s_{1}},\cdots,A_{s_{m}})$ and $(A_{s_{1}}^{\prime},\cdots,A_{s_{m}}^{\prime})$ over the same domains, we define
@@
4.3. First Attempt at Defining Judgments
This section presents a first attempt to introduce a framework for quantitative local reasoning in general quantum programs using tuples of local observables.
In Quantum Hoare Logic (QHL) (Ying11, ), the proof rule for a unitary operation $P$ takes the form
$\{U_{P}^{\dagger}BU_{P}\}\ P\ \{B\},$
which characterizes the weakest precondition for the postcondition $B$.
Let $S=(s_{1},\ldots,s_{m})$ be an $m$-tuple with each $s_{i}\subseteq[n]$, and consider a predicate $\mathscr{B}=(B_{s_{1}},\ldots,B_{s_{m}})$; we have
$M_{\mathscr{B}}=\sum_{i}B_{s_{i}}\otimes I_{[n]\setminus s_{i}},$
where $0\leq B_{s_{i}}\leq I_{s_{i}}$.

### v3

@@ 2.2. Quantum States
[…]
@@ 2.3. Reduced Density Matrices
[…]
For an $n$-partite system and $s\subseteq[n]$, the reduced density matrix is
@@
with ${\mathrm{Tr}}_{[n]\setminus s}$ tracing out all qubits not in $s$. The partial trace preserves positive semi-definiteness (NI11, ).
@@ 2.4. Unitary Operations
[…]
@@ 2.5. Observables
Quantum observables represent measurable quantities, such as position, energy, or spin, and are modeled by Hermitian operators $O$ with $O^{\dagger}=O$, ensuring real outcomes. In contexts like Quantum Hoare Logic (QHL), observables are often restricted to $0\leq O\leq I$ in the Löwner order (§˜2.1), meaning both $O$ and IOI-O are positive semidefinite. These bounded observables serve as predicates or quantum effects, capturing partial truth values in program verification.
@@ 2.6. Quantum Circuits and Semantics
[…]
Each gate is lifted to an $n$-qubit unitary by tensoring with identities. For a single-qubit gate $U$ acting on qubit $i$, $U\otimes I_{[n]\setminus\{i\}}:=(\otimes_{k>i}I)\otimes U\otimes(\otimes_{0<j<i}I)$, and similarly for two-qubit gates $U_{F_{\ell}}$ on $F_{\ell}=\{i,j\}$, $U\otimes I_{[n]\setminus\{i,j\}}$, with placement determined by the qubit indices. For clarity, we describe semantics assuming 2-qubit gates, though all results extend to gates acting on up to $m$ qubits for any constant $m$.
@@
We also define the dual action on observable $A$ as:
$\llbracket C\rrbracket^{*}(A):=U_{C}^{\dagger}AU_{C}.$
We define $\mathcal{U}:=\lambda x.UxU^{\dagger}$ to denote the quantum operation on reduced density matrices that corresponds to the unitary matrix $U$ (which operates on quantum states).
@@ 2.7. Projections
An orthogonal projection satisfies $P=P^{\dagger}=P^{2}$, a stronger condition than the classical $P=P^{2}$. We simply call such matrices projections. Each projection $P$ corresponds to a subspace $S_{P}=\{v\mid Pv=v\}$, and we use “projection” and “subspace” interchangeably. This induces a partial order: $P\subseteq Q$ iff $S_{P}\subseteq S_{Q}$. For example, $|00\rangle\langle 00|+|11\rangle\langle 11|$ is a rank-2 projection onto a 2D subspace of $\mathbb{C}^{4}$.
Projections are positive semidefinite. The support of a positive semidefinite matrix $A$, $\operatorname{supp}(A)$, is the span of eigenvectors with nonzero eigenvalues. A density matrix $\rho$ satisfies $P$, written $\rho\vDash P$, if $\operatorname{supp}(\rho)\subseteq P$, equivalently $P\rho=\rho$ (BvN36, ).
@@ 2.8. Lemmas
The following lemmas summarize key algebraic relationships among these operations, forming the basis for the correctness of our SAQR-QC framework. Unless stated otherwise, let $s\subseteq[n]$; $P$ be a projection on an $n$-qubit system; $A,B$ be positive semidefinite matrices; $E$ a matrix; and $\rho$ a quantum state. The proof of ˜2.1 is given in Appendix˜A. The other lemmas can be proven using the definition of trace, partial trace, and support.
The relevance of these lemmas to SAQR-QC is that (i) a positive semidefinite matrix $A$ serves as a predicate on a density matrix $\rho$ via the expectation value ${\mathrm{Tr}}(A\rho)$; and (ii) a subclass of such predicates with good locality properties are ones defined by a positive semidefinite matrix $A_{s}$ that only acts on qubit set $s\subseteq[n]$, producing the expectation value ${\mathrm{Tr}}\left((A_{s}\otimes I_{[n]\setminus s})\rho\right)={\mathrm{Tr}}\left(A_{s}\rho_{s}\right)$.
@@
Let $\rho$ be the density matrix of an $n$-qubit system, and let $s\subseteq[n]$. Then for any observable $A_{s}$ acting on subsystem $s$, ${\mathrm{Tr}}\left((A_{s}\otimes I_{[n]\setminus s})\rho\right)={\mathrm{Tr}}\left(A_{s}\rho_{s}\right)$.
@@
If $A\leq B$, then ${\mathrm{Tr}}(A\rho)\leq{\mathrm{Tr}}(B\rho)$ for any density operator $\rho$, and in particular ${\mathrm{Tr}}(A)\leq{\mathrm{Tr}}(B)$.
@@
2.9. Qualitative Predicates for Local Reasoning: Quantum Abstract Interpretation
The qualitative predicates used in QAI express logical properties of quantum states, e.g., whether a subsystem lies in a subspace. Represented as tuples of local projectors, they support scalable local reasoning, inferring global behavior from partial views. QAI (YP21, ) propagates these predicates through circuits using small-subsystem projectors, avoiding exponential full-state analysis.
This section reviews local projective predicates, formalizes their semantics, and describes their transformation under unitaries. Soundness is established via support-based semantics and partial trace, forming the basis for abstract reasoning about quantum programs.
@@ 2.9. Qualitative Predicates for Local Reasoning: Quantum Abstract Interpretation
A state $\rho$ satisfies a projective predicate $\mathcal{P}=(P_{s_{1}},\dots,P_{s_{m}})$, denoted by
@@
if for all $1\leq i\leq m$, $P_{s_{i}}\rho_{s_{i}}=\rho_{s_{i}}$, i.e., $\rho_{s_{i}}\vDash P_{s_{i}}$. Equivalently, $\rho\vDash\gamma(\mathcal{P})$, where
@@
[…]
@@ 3. Correctness Formulas and the Logical System SAQR-QC
In this section, we introduce a class of predicates that can be used for quantitative local reasoning (§˜3.1). We then sketch a strawman approach that incorporates these predicates in a logic similar to Quantum Hoare Logic (QHL) (§˜3.2). Via a simple example, we demonstrate that this strawman approach can lead to a significant loss of precision in reasoning. To address this limitation, we integrate the QAI technique into the framework, leading to the formal definition of the judgments used in SAQR-QC (§˜3.3), a theorem about how SAQR-QC judgments relate to judgments in QAI (§˜3.4), a systematic presentation of the inference rules of SAQR-QC (§˜3.5), and an approximation strategy that addresses how to work with the one inference rule of SAQR-QC that poses a challenge to scalability (§˜3.6).
@@ 3.1. Quantitative Local Reasoning via Generalized Predicates
Reasoning about full quantum states quickly becomes infeasible due to the exponential growth of the state space with the number of qubits. To address this issue, we want an approach akin to QAI (YP21, ), which uses tuples of projections as predicates that capture properties of small subsystems. However, QAI itself is unsatisfactory because QAI can only perform qualitative local reasoning.
To enable quantitative local reasoning, we focus on reduced density matrices, which capture the behavior of local subsystems. Inspired by Quantum Hoare Logic (QHL)—where a global positive semidefinite matrix $A$ serves as a predicate describing properties of a state $\rho$ via the expectation value ${\mathrm{Tr}}(A\rho)$—we introduce the notion of local observables to reason about subsystem states $\rho_{s}$.
@@
A local observable over a qubit set $s\subseteq[n]$ is a positive semidefinite operator $A_{s}$ satisfying $0\leq A_{s}\leq I_{s}$, acting nontrivially only on the qubits in $s$.
According to Lemma 2.1, ${\mathrm{Tr}}(A_{s}\rho_{s})={\mathrm{Tr}}(A\rho)$ with $A:=A_{s}\otimes I_{[n]\setminus s}$. Local observables thus form a natural subclass of predicates in QHL, and we use them to monitor reduced density matrices via the inner product ${\mathrm{Tr}}(\cdot)$, in direct analogy with QHL; here, “monitoring $\rho_{s}$” means computing the expectation value of local observable $A_{s}$ with respect to $\rho_{s}$—i.e., computing ${\mathrm{Tr}}(A_{s}\rho_{s})$.
To capture richer information, we track not a single $\rho_{s}$ but a tuple of reduced density matrices $(\rho_{s_{1}},\dots,\rho_{s_{m}})$, where each $s_{i}\subseteq[n]$ is a small subset of bounded size; we also refer to these as low-dimensional marginals. This representation (i) provides significantly deeper insight into the system’s local structure, and (ii) remains tractable, because it requires only linear resources.
@@
A predicate over an $n$-qubit space and $S=(s_{1},\dots,s_{m})$ is a tuple of local observables
$(A_{s_{1}},\dots,A_{s_{m}})\quad\text{with}\quad 0\leq A_{s_{i}}\leq I_{s_{i}},$
each acting nontrivially only on $s_{i}$. Predicates are denoted by $\mathscr{A}$ (or $\mathscr{B},\mathscr{D}$); $m$ is the predicate’s size.
Each local observable acts nontrivally on just a few qubits, and captures the expectation value of a measurable quantity, such as the success probability, fidelity, or entanglement, providing a direct link between semantics and operational outcomes. By tracing their evolution under unitaries and partial traces, we obtain a scalable, compositional framework for quantum program verification, supporting precise analysis without reconstructing the full state.
These generalized predicates track quantitative aspects of quantum programs—such as probabilities and expectation values—by monitoring reduced density matrices over small, constant-size subsystems. This abstraction underpins scalable, compositional reasoning about quantum computations via low-dimensional, information-preserving summaries.
Definition 3.3.
@@
$\small M_{\mathscr{A}}=\sum_{i=1}^{m}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}.$
This definition tracks changes in a tuple of reduced density matrices via linear functionals, enabling scalable, compositional quantitative reasoning. By Lemma 2.1, the trace ${\mathrm{Tr}}(M_{\mathscr{A}}\rho)$ depends only on the reduced states $\rho_{s_{i}}$:
(4) $\small{\mathrm{Tr}}(M_{\mathscr{A}}\rho)={\mathrm{Tr}}\Big(\sum_{i=1}^{m}(A_{s_{i}}\otimes I_{[n]\setminus s_{i}})\rho\Big)=\sum_{i=1}^{m}{\mathrm{Tr}}[(A_{s_{i}}\otimes I_{[n]\setminus s_{i}})\rho]=\sum_{i=1}^{m}{\mathrm{Tr}}(A_{s_{i}}\rho_{s_{i}}).$
@@
The domain of a predicate $(A_{s_{1}},\dots,A_{s_{m}})$ is $\mathrm{dom}(A_{s_{1}},\dots,A_{s_{m}})=(s_{1},\dots,s_{m})$.
@@
Given predicates $(A_{s_{1}},\cdots,A_{s_{m}})$ and $(A_{s_{1}}^{\prime},\cdots,A_{s_{m}}^{\prime})$ over the same domains, define
@@
3.2. A First Attempt at Defining Judgments
This section presents an initial attempt for quantitative local reasoning in quantum programs using tuples of local observables.
In QHL (Ying11, ), the unitary rule gives the weakest precondition for a postcondition:
$\{U_{P}^{\dagger}BU_{P}\}\ P\ \{B\}.$
Let $S=(s_{1},\dots,s_{m})$ with $s_{i}\subseteq[n]$, and $\mathscr{B}=(B_{s_{1}},\dots,B_{s_{m}})$ with $0\leq B_{s_{i}}\leq I_{s_{i}}$; then
$M_{\mathscr{B}}=\sum_{i}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}.$